OpenStudy (anonymous):

4/3 *t=-5/2???? help

OpenStudy (anonymous):

you mean \[(4/3)t=-5/2\] or \[4/(3t)=-5/2\]

OpenStudy (anonymous):

the first

OpenStudy (anonymous):

okk then multiply both sides by 3/4

OpenStudy (anonymous):

to get rid of (4/3)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

you want a full answer ? :) ?

OpenStudy (anonymous):

well i need help to understand what to do

OpenStudy (anonymous):

so it would be 15/8

OpenStudy (anonymous):

so close :) -15/8 :)

OpenStudy (anonymous):

can u explain to me the operations with neg and post numbers i ddont know how it goes

OpenStudy (anonymous):

i kno two neg makes a pos

OpenStudy (anonymous):

\[(4/3)t=-5/2\] we need to "isolate" t so what should we do ! multply both sides by 3/4 because (3/4)*(4/3)t=t did you get or not :) !

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry ! I'm not a good teacher :( !

OpenStudy (anonymous):

oh u did good fractions are hard for me

OpenStudy (anonymous):

yeaah ! I know...It was hard for me too ; one day ! just practise more ! :)

OpenStudy (anonymous):

\[-18= -20a\]

OpenStudy (anonymous):

do i multiply by the -20

OpenStudy (anonymous):

no ; we had 4/3 and we reverse it we multiplied b 3/4 now !reverse -20 what would you get ?! :)

OpenStudy (anonymous):

1/20

OpenStudy (anonymous):

and u didn't forget anything ?! :)

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

gooood ! the answer is good ! but you had to multiply by -1/20 ! I guess this is what you di :) :)

OpenStudy (anonymous):

did*

OpenStudy (anonymous):

yeea so the answer is -4/5

OpenStudy (anonymous):

just did one by myself tell me if i got it

OpenStudy (anonymous):

okk if you (-1/20)*(-18) negative by negative = ?

OpenStudy (anonymous):

\[4/5\div8=4/5 *1/8=4/40=1/10\]

OpenStudy (anonymous):

yeaaaah !! good !

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

yeaaah so the correct answer was ... ?!

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

yeaaah !! that's correct !

OpenStudy (anonymous):

how many 2/3 cups cereal bowl frow 10 cups of cereal\[2/3 *1/10= 2/30=1/15\]

OpenStudy (anonymous):

from*

OpenStudy (anonymous):

it can't be true ! why ?! because 1/15 is not even one cup ! you can correct it !

OpenStudy (anonymous):

were did i go wrong

OpenStudy (anonymous):

(2/3) *10

OpenStudy (anonymous):

did you get it ! ?!

OpenStudy (anonymous):

i se e my mistake

OpenStudy (anonymous):

is it 15

OpenStudy (anonymous):

not really :)! Sorry ; :( I got to go ! :(

OpenStudy (anonymous):

k